Carbon disulfide is produced by the reaction of carbon and sulfur dioxide. What is the percent yield of carbon disulfide if the reaction of 180.0 g of sulfur dioxide produces 72.0 grams of carbon disulfide? (3 pts.)

                               5C  +  2SO2 --> CS2  + 4CO

Show work


1
Expert's answer
2021-07-09T05:15:21-0400

Moles of SO2 reacted = mass / molar mass

= 32.0/ 64.0

= 0.5 mol


2 moles of SO2 forms 1 mol of CS2

so,

theoretical moles of CS2 = 0.5/2 = 0.25 mol


theroetical mass of CS2 = number of mol ×molar mass

= 0.25 mol ×76.1 g/mol

= 19.0 g


actual mass = 10.2 g

percentage yield = actual mass × 100 / theroetical mass

= 10.2*100 / 19.0

=53.7 %

Answer: 53.7 %